# publications.bib

### v3

@@ 4.1. The Spectral Action
Now, we want to define $\mathrm{LLC}_{\mathrm{Bun}_{\mathrm{GU}_{2}}}$ in terms of $\mathrm{LLC}_{\mathrm{Bun}_{H}}$. More precisely, for any $b^{\prime}\in B(\mathrm{GU}_{2})$ we define $\mathrm{LLC}_{b^{\prime}}:\Pi(J_{b^{\prime}})\rightarrow\Phi(J_{b^{\prime}})$ in terms of $\mathrm{LLC}_{b}:\Pi(J_{b})\rightarrow\Phi(J_{b})$ for $\mathrm{LLC}_{\mathrm{Bun}_{H}}$. For $\pi_{b^{\prime}}=(\tilde{\pi},\chi)\in\Pi(J_{b^{\prime}})$, we consider the image $\phi=\mathrm{LLC}_{\mathrm{Bun}_{H}}((\tilde{\pi},\chi))$, and we want to show that $\phi:\mathrm{WD}_{L}\rightarrow\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})$ factors through $\phantom{}{}^{L}J_{b^{\prime}}(\overline{\mathbb{Q}}_{\ell})$. To see this, from the exact sequence above, it suffices to show that the composition of $\phi$ with the map $\phantom{}{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}\mathbb{G}_{m}(\overline{\mathbb{Q}}_{\ell})$ is trivial. However, we observe that the condition that $\chi|_{L^{\times}}=\omega_{\tilde{\pi}}|_{L^{\times}}$ exactly implies that this image is trivial, since the composition is the $L$-parameter associated with the character $\omega_{\tilde{\pi}}\chi^{-1}|_{L^{\times}}$. Thus, we have an $L$-parameter $\phi^{\prime}:\mathrm{WD}_{L}\rightarrow\phantom{}^{L}J_{b^{\prime}}(\overline{\mathbb{Q}}_{\ell})$. We thus define the map $\mathrm{LLC}_{b^{\prime}}$ to take $\pi_{b^{\prime}}$ to $\phi^{\prime}$.
We now have the following theorem about the groups we know to satisfy Assumption 4.4.
@@
[Ham1, FS, HKW, BMNH] Assumption 4.4 is true and $\ell$ is very good in the following cases.
@@
    The group $\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GSp}_{4})$ with $p>2$ and $[L:\mathbb{Q}_{p}]\geq 2$ or $L=\mathbb{Q}_{p}$ for all $p$. In both cases, we need to assume that $\ell\nmid 2[L:\mathbb{Q}_{p}]$.
@@
    The group $\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GU}_{2})$ defined with respect to an unramified quadratic extension $L^{\prime}/L$, and $\ell$ such that $\ell\nmid[L:\mathbb{Q}_{p}]$.
@@
We first start with the conditions on $\ell$. If $G$ is of type $A_{n}$ then all $\ell$ are very good. However, when $G$ is a unitary group with $n>2$, we also need to impose the additional assumption that the action of $W_{\mathbb{Q}_{p}}$ on $\hat{G}$ is of order prime to $\ell$, and this gives us the condition that $\ell\neq 2$. Similarly, this gives rise to the condition that $\ell\nmid[L:\mathbb{Q}_{p}]$ in all of the cases. If $G=\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GSp}_{4})$ then it is of type $C$ and we also need to impose the additional condition that $\ell\neq 2$.
@@
Now, we turn to Assumption 4.4 (1). For $\mathrm{GL}_{n}$, this follows from [FS, Theorem I.9.6] and [HKW, Theorem 1.0.3], where $\mathrm{LLC}_{b}$ is given by the Harris-Taylor correspondence. For $\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GSp}_{4})$ and $L/\mathbb{Q}_{p}$ as described above, this follows from [Ham1, Theorem 1.1], where $\mathrm{LLC}_{b}$ is given by Harris-Taylor for the non-basic $b$ and Gan-Takeda [GT1] and Gan-Tantono [GT2] for the basic elements [^6] In the current version of [Ham1], the assumption that $p>2$ is only used to invoke basic uniformization of abelian type Shimura varieties, but when $L=\mathbb{Q}_{p}$ one can just use Rappoport-Zink uniformization, so this assumption is unnecessary.. For $\mathrm{U}_{n}$ or $\mathrm{GU}_{n}$, this is [BMNH, Theorem 1.1], where $\mathrm{LLC}_{b}$ for $b\in B(G)$ was constructed by Mok [Mok] and Kaletha-Minguez-Shin-White [KMSW]. For $\mathrm{GU}_{2}$ this follows from the compatibility for $\mathrm{GL}_{2}$, and the fact that the Fargues-Scholze local Langlands correspondence is compatible with taking products as well as maps $G^{\prime}\rightarrow G$ that induces an isomorphism of adjoint groups [FS, Theorem I.9.6 (v),(vi)].
@@
Now we explain why Assumption 4.4 (2) is satisfied. We recall that if $J_{b}$ is a non quasi-split group then the fibers of the $\mathrm{LLC}_{b}$ over an $L$-parameter $\phi:\mathrm{WD}_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ should be empty if $\phi$ factors through $\phantom{}{}^{L}M(\overline{\mathbb{Q}}_{\ell})$ for a Levi subgroup $M\subset G$ which does not transfer to a Levi subgroup of $J_{b}$. In particular, such parameters are called irrelevant, and we expect the fiber to be empty if and only if $\phi$ is irrelevant [Kal, Conjecture A.2]. For the Harris-Taylor correspondence, it is know that the fibers over irrelevant parameters are empty by the standard properties of Jacquet-Langlands. For $\mathrm{GU}_{n}$ or $\mathrm{U}_{n}$, odd unitary groups and their Levi subgroups are always quasi-split, so it is reduced to the previous case of $\mathrm{GL}_{n}$ using compatibility of the correspondence with parabolic induction. For $\mathrm{GSp}_{4}$, one needs to show this for $\mathrm{LLC}_{\mathrm{GU}_{2}(D)}$, where $\mathrm{GU}_{2}(D)$ is the unique non-split inner form of $\mathrm{GSp}_{4}$. Here this follows from the construction of Gan-Tantono (See the discussion before the main Theorem in [GT2]). For $\mathrm{GU}_{2}$, we observe that since $\mathrm{GU}_{2},H,\mathrm{GL}_{2}$ all have the same adjoint group, $b^{\prime}\in B(\mathrm{GU}_{2})_{\mathrm{un}}$ is unramified exactly when $b$ is unramified with notation as in Lemma 4.9. Now, let $\tilde{T}$ be a maximal split torus of $\mathrm{GL}_{2}$, and observe that $T^{\prime}=(\tilde{T}\times\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m})/\mathbb{G}_{m}$ is a maximal torus of $\mathrm{GU}_{2}$ over $L$. Given $\pi_{b^{\prime}}=(\tilde{\pi},\chi)$, and $\phi_{\pi_{b^{\prime}}}:\mathrm{WD}_{L}\rightarrow\phantom{}^{L}J_{b^{\prime}}(\overline{\mathbb{Q}}_{\ell})$, we see from the construction that this factors through $\phantom{}{}^{L}T^{\prime}(\overline{\mathbb{Q}}_{\ell})$ exactly when the associated $L$-parameter for $H$, $\phi_{(\tilde{\pi},\chi)}:\mathrm{WD}_{L}\rightarrow\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})$, factors through $\phantom{}{}^{L}T(\overline{\mathbb{Q}}_{\ell})$, where $T=\tilde{T}\times\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m}$. Since Assumption 4.4 (2) is clearly compatible with taking products, $H$ satisfies this assumption, and thus so does $\mathrm{GU}_{2}$.
@@
Now we explain why Assumption 4.4 (3) is satisfied. First, note that any parameter $\phi:\mathrm{WD}_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ induced from a toral parameter $\phi_{T}$ has necessarily trivial monodromy, since $\phantom{}{}^{L}T(\overline{\mathbb{Q}}_{\ell})$ consists only of semi-simple elements. Moreover, since $J_{b}$ is an inner form of $M_{b}$, it follows that the set of all distinct conjugacy classes of parameters $\phi^{\prime}:\mathrm{WD}_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})$ which can give rise to $\phi$ under the twisted embedding $\phantom{}{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ are parameterized by a set of minimal length representatives of $W_{b}=W_{G}/W_{M_{b}}$ via conjugating $\phi^{\prime}$. We expect (See [Kal, Conjecture A.5]) that the fiber of $\mathrm{LLC}_{b}$ over such a $\phi^{\prime}$ inducing $\phi$ to be the irreducible constituents of the normalized induction of the $L$-packet of $\phi_{T}^{w}$, which is just $\chi^{w}$ by local class field theory for $w\in W_{b}$. This is indeed true in all the cases we consider (See for example [BMNH, Section 2.3.3] for this discussed in the case of unitary groups, and for $\mathrm{GSp}_{4}$ and its unique non-split inner form $\mathrm{GU}_{2}(D)$ it follows directly from the construction). We note that the twists by $\delta_{P_{b}}^{-1/2}$ appear to account for the half Tate twists appearing in the definition of the twisted embedding $\phantom{}{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$. For $\mathrm{GU}_{2}$, we see that when $b^{\prime}$ is unramified, we have isomorphisms of flag varieties
@@
We now turn our attention to deriving our desired consequences.

### v5

@@ 4.2.1. Implications of Compatibility for Generic Parameters
Now, we want to define $\mathrm{LLC}_{\mathrm{Bun}_{\mathrm{GU}_{2}}}$ in terms of $\mathrm{LLC}_{\mathrm{Bun}_{H}}$. More precisely, for any $b^{\prime}\in B(\mathrm{GU}_{2})$ we define $\mathrm{LLC}_{b^{\prime}}:\Pi(J_{b^{\prime}})\rightarrow\Phi(J_{b^{\prime}})$ in terms of $\mathrm{LLC}_{b}:\Pi(J_{b})\rightarrow\Phi(J_{b})$ for $\mathrm{LLC}_{\mathrm{Bun}_{H}}$. For $\pi_{b^{\prime}}=(\tilde{\pi},\chi)\in\Pi(J_{b^{\prime}})$, we consider the image $\phi=\mathrm{LLC}_{\mathrm{Bun}_{H}}((\tilde{\pi},\chi))$, and we want to show that $\phi:W_{L}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})$ factors through $\phantom{}{}^{L}J_{b^{\prime}}(\overline{\mathbb{Q}}_{\ell})$. To see this, from the exact sequence above, it suffices to show that the composition of $\phi$ with the map $\phantom{}{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}\mathbb{G}_{m}(\overline{\mathbb{Q}}_{\ell})$ is trivial. However, we observe that the condition that $\chi|_{L^{\times}}=\omega_{\tilde{\pi}}|_{L^{\times}}$ exactly implies that this image is trivial, since the composition is the $L$-parameter associated with the character $\omega_{\tilde{\pi}}\chi^{-1}|_{L^{\times}}$. Thus, we have an $L$-parameter $\phi^{\prime}:W_{L}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b^{\prime}}(\overline{\mathbb{Q}}_{\ell})$. We thus define the map $\mathrm{LLC}_{b^{\prime}}$ to take $\pi_{b^{\prime}}$ to $\phi^{\prime}$.
For the various groups discussed above, we will also want to check some assumptions on the prime $\ell$ relative to the group $G$ that we will need to assume for the proof our main results, since they are required for the arguments in [24]. In particular, we define the following.
Definition 4.11.
We say that $\ell$ is very decent with respect to $G$ if the following holds.
  1. (1)
    We have that $\ell\nmid|\pi_{0}(Z(G))|$, where $Z(G)$ denotes the center of $G$.
  2. (2)
    We have that $\ell\nmid|Q|$, where we recall that $Q$ is the smallest quotient through which the action of $W_{\mathbb{Q}_{p}}$ on $\hat{G}$ factors.
Remark 4.12.
The motivation for the first condition was already explained in §4.1; in particular, it is required for the construction of the spectral action and in turn the Fargues-Scholze local Langlands correspondence and the localized derived category $\mathrm{D}(\mathrm{Bun}_{G},\overline{\mathbb{F}}_{\ell})_{\phi}$. The second assumption comes from results on the theory of geometric Eisenstein series introduced in [24]. In particular, it will guarantee that the theory of tilting modules for the group $\hat{G}/\overline{\mathbb{F}}_{\ell}$ extends to a nice theory of tilting modules for the group $\phantom{}{}^{L}G:=\hat{G}\ltimes Q$, as discussed in [24, Section 9], which is used to establish the main results of [24]. In particular, the semisimplicity of the category of tilting modules for the full $L$-group $\phantom{}{}^{L}G$ requires this.
We now have the following Theorem about the groups we know to satisfy Assumption 4.5, preempting our discussion in the next section.
@@
[27, 16, 31, 47] Assumption 4.5 is true and $\ell$ is very decent in the following cases.
@@
    The group $\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GSp}_{4})$ for $L/\mathbb{Q}_{p}$ an unramified extension with $p>2$ , $L:\mathbb{Q}_{p}]\geq 2$, and $\ell\nmid[L:\mathbb{Q}_{p}]$ or $L=\mathbb{Q}_{p}$ for all $p$ and $\ell$.
@@
    The group $\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GU}_{2})$ defined with respect to an unramified quadratic extension $L^{\prime}/L$, and $\ell$ such that $\ell\nmid 2[L:\mathbb{Q}_{p}]$.
@@
We first start with the conditions on $\ell$. For all the groups described above the center is connected, and therefore condition (1) of Definition 4.11 is automatically satisfied. For condition (2) of Definition 4.11, it is straightforward. In particular, in the unramified restriction of scalar cases discussed in (1) and (4), the quotient $Q$ will be isomorphic to the Galois group of $L/\mathbb{Q}_{p}$, and in the unitary group case in (2) it will be isomorphic to the Galois group of the quadratic extension $E/\mathbb{Q}_{p}$. In case (3), it will be given by the Galois group of the extension $L^{\prime}/\mathbb{Q}_{p}$.
@@
Now, we turn to Assumption 4.5 (1). For $\mathrm{GL}_{n}$, this follows from [16, Theorem I.9.6] and [31, Theorem 1.0.3], where $\mathrm{LLC}_{b}$ is given by the Harris-Taylor correspondence precomposed with Badulescu’s Jacquet Langlands map [2]. For $\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GSp}_{4})$ and $L/\mathbb{Q}_{p}$ as described above, this follows from [27, Theorem 1.1], where $\mathrm{LLC}_{b}$ is given by Harris-Taylor for the non-basic $b$ and Gan-Takeda [19] and Gan-Tantono [20] for the basic elements [^10] In the current version of [27], the assumption that $p>2$ is only used to invoke basic uniformization of abelian type Shimura varieties, but this condition is actually unnecessary if $L=\mathbb{Q}_{p}$, following the construction of Pappas-Rapoport [50] of the uniformization in Hodge type case when $p=2$ and the group is unramified.. For $\mathrm{U}_{n}$ or $\mathrm{GU}_{n}$, this is [47, Theorem 1.1], where $\mathrm{LLC}_{b}$ for $b\in B(G)$ was constructed by Mok [48] and Kaletha-Minguez-Shin-White [34].
@@
Now we explain why Assumption 4.5 (2) is satisfied. We recall that if $J_{b}$ is a non quasi-split group then the fibers of the $\mathrm{LLC}_{b}$ over an $L$-parameter $\phi:W_{\mathbb{Q}_{p}}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})$ should be empty if $\phi$ factors through $\phantom{}{}^{L}M(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ via the the natural (equivalently twisted) embedding for a Levi subgroup $M\subset G$ which does not transfer to a Levi subgroup of $J_{b}$ (recalling that $G$ is quasi-split in this case, so that $J_{b}$ is an inner form of the quasi-split Levi of $G$). In particular, such parameters are called irrelevant, and we expect the fiber to be empty if and only if $\phi$ is irrelevant [35, Conjecture A.2]. Assumption 4.5 (2) reduces to checking that the fibers over the irrelevant $L$-parameters are empty for all $J_{b}$ and the $\mathrm{LLC}_{b}$ in question. This is because $b\in B(G)_{\mathrm{un}}$ if and only if $J_{b}$ is quasi-split or a torus (See [24, Lemma 2.12]) implying that if $b\notin B(G)_{\mathrm{un}}$ then all $L$-parameters $\phi\in\Phi(J_{b})$ which factor through $\phantom{}{}^{L}T$ are irrelevant. Since every representation in $\Pi(J_{b})$ occurs as the constituent of a parabolic induction of a supercuspidal representation of a Levi subgroup (where we recall that supercuspidals are in particular discrete series) to show this claim, it suffices to show that the semi-simplification of $\mathrm{LLC}_{b}$ is compatible with respect to parabolic induction and the semi-simplification of a local Langlands correspondence on each Levi $L_{b}\subset J_{b}$ (as in [16, Theorem I.9.6 (viii)]), and that the non-semisimplified local Langlands correspondence on the Levi sends discrete series representations to discrete $L$-parameters of $L_{b}$ (in the sense that they do not factor through some smaller Levi of $\phantom{}{}^{L}L_{b}$), as expected [35, Conjecture A.4]).
For the Harris-Taylor correspondence, it is know that the fibers over irrelevant parameters are empty by the standard properties of Jacquet-Langlands. In particular, it follows from the discussion before [40, Theorem 4.2.1] that the Harris-Taylor parameter for a discrete series representation of $\mathrm{GL}_{n}$ always defines a discrete L-parameter (in the sense that its L-parameter does not factor through a proper Levi subgroup) and is compatible with parabolic induction. The claim for $\mathrm{GL}_{m}(D_{\frac{1}{n}})$ now follows from the fact that the Jacquet Langlands transfer of any smooth irreducible representation of $D_{\frac{1}{n}}^{*}$ is discrete, and the compatibility of parabolic induction of Baduluescu’s Jacquet Langlands transfer discussed in [2, Section 3.1]. This implies the claim on irrelevant parameters for $J_{b}$ in the case of $G=\mathrm{GL}_{n}$, since $J_{b}$ in this case will always be a product of groups of the form $\mathrm{GL}_{m}(D_{\frac{1}{n}})$.
For $G=\mathrm{GU}_{n}$ or $\mathrm{U}_{n}$, we note that odd unitary groups are always quasi-split, so that the inner forms of such a group are always quasi-split; in particular, there are no irrelevant $L$-parameters for these groups and there is nothing to check. Therefore in this case, we can reduce to the claim on irrelevant $L$-parameters to $\mathrm{GL}_{n}$ and its inner forms, since these occur as direct factors of $J_{b}$ for such $G$, which was already explained above.
For $\mathrm{GSp}_{4}$, one needs to show the claim on irrelevant $L$-parameters for $\mathrm{LLC}_{\mathrm{GU}_{2}(D)}$, where $\mathrm{GU}_{2}(D)$ is the unique non-split inner form of $\mathrm{GSp}_{4}$. Here the claim on irrelevant $L$-parameters follows directly from the construction of Gan-Tantono (See the discussion before the main Theorem in [20]).
For $\mathrm{GU}_{2}$, we observe that since $\mathrm{GU}_{2}$$H:=\mathrm{GL}_{2}\times\mathrm{Res}_{L^{\prime}/L}(\mathbb{G}_{m})$ and, $\mathrm{GL}_{2}$ all have the same adjoint group, $b^{\prime}\in B(\mathrm{GU}_{2})_{\mathrm{un}}$ is unramified exactly when $b$ is unramified with notation as in Lemma 4.10. Now, let $\tilde{T}$ be a maximal split torus of $\mathrm{GL}_{2}$, and observe that $T^{\prime}=(\tilde{T}\times\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m})/\mathbb{G}_{m}$ is a maximal torus of $\mathrm{GU}_{2}$ over $L$. Given $\pi_{b^{\prime}}=(\tilde{\pi},\chi)$, and $\phi_{\pi_{b^{\prime}}}:W_{L}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b^{\prime}}(\overline{\mathbb{Q}}_{\ell})$, we see from the construction that this factors through $\phantom{}{}^{L}T^{\prime}(\overline{\mathbb{Q}}_{\ell})$ exactly when the associated $L$-parameter for $H$, $\phi_{(\tilde{\pi},\chi)}:W_{L}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})$, factors through $\phantom{}{}^{L}T(\overline{\mathbb{Q}}_{\ell})$, where $T=\tilde{T}\times\mathrm{Res}_{L^{\prime}/L}\mathbb{G}_{m}$. Since Assumption 4.5 (2) is clearly compatible with taking products, the group $H$ via $a\mapsto(\mathrm{diag}(a,a),a^{-1})$ satisfies this assumption. Thus, we see that so does $\mathrm{GU}_{2}$, since if we let $T_{H}$ denote the maximal torus of $T$ such that its image under the projection to $\mathrm{GU}_{2}$ is $T$, and we let $\rho_{H}$ be a smooth irreducible representation of $H(L)$ lifting $\rho$, we see that by definition of $\mathrm{LLC}_{\mathrm{GU}_{2}}$ we have $\phi_{\rho}$ factors through $\phantom{}{}^{L}T$ if and only if $\phi_{\rho_{H}}$ factors through $\phantom{}{}^{L}T_{H}$.
@@
Now we explain why Assumption 4.5 (3) is satisfied. First, note that any parameter $\phi:W_{\mathbb{Q}_{p}}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ induced from a toral parameter $\phi_{T}$ has necessarily trivial monodromy, since $\phantom{}{}^{L}T(\overline{\mathbb{Q}}_{\ell})$ consists only of semi-simple elements. Moreover, since $J_{b}$ is an inner form of $M_{b}$, it follows that the set of all distinct conjugacy classes of parameters $\phi^{\prime}:W_{\mathbb{Q}_{p}}\times\mathrm{SL}_{2}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})$, which can give rise to $\phi$ under the twisted embedding $\phantom{}{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ are parameterized by a set of minimal length representatives of $W_{b}=W_{G}/W_{M_{b}}$ via conjugating $\phi^{\prime}$. We expect (See [35, Conjecture A.5]) that the fiber of $\mathrm{LLC}_{b}$ over such a $\phi^{\prime}$ inducing $\phi$ to be the irreducible constituents of the normalized induction of the $L$-packet of $\phi_{T}^{w}$, which is just $\chi^{w}$ by local class field theory for $w\in W_{b}$. For cases (1),(2),and (4), the required compatibility with parabolic induction above was already discussed above for the Harris-Taylor correspondence and the relevant claim for the Kaletha-Minguez-Shin-White/Mok correspondence follows from [47, Section 2.2.3]. For the Gan-Takeda/Gan-Tantono correspondence, it follows directly from the construction and the analogous claim for the Harris-Taylor correspondence. We note that the twists by $\delta_{P_{b}}^{-1/2}$ appear to account for the half Tate twists appearing in the definition of the twisted embedding $\phantom{}{}^{L}J_{b}(\overline{\mathbb{Q}}_{\ell})\rightarrow\phantom{}^{L}G(\overline{\mathbb{Q}}_{\ell})$ (cf. [25, Another proof of Proposition 4.1]). For $\mathrm{GU}_{2}$, we see that when $b^{\prime}$ is unramified, we have isomorphisms of flag varieties
@@
We now turn our attention to deriving the key local results from Assumption 4.5.